OpenStudy (anonymous):

lim x tanx/ 1-cosx x―>0 please help...

OpenStudy (anonymous):

Try multiplying the numerator and denominator by 1+cosx. See if you can solve it using that.

OpenStudy (aravindg):

xtan x(1+cos x)/1-cos x)(1+cos x)

OpenStudy (aravindg):

then xtanx/sin^2x

OpenStudy (aravindg):

div by x^2 in num and denominator

OpenStudy (anonymous):

(You forgot the 1+cosx factor in the numerator, AvarindG)

OpenStudy (aravindg):

srry ringlum

OpenStudy (aravindg):

wel katy i suggest u work out rest and a[ply limits .u will get it

OpenStudy (anonymous):

okie....thank u both...i got stuck wid the div part.. :)

OpenStudy (anonymous):

It's ok. In general for limits involving trigonometric functions, keep in mind the identities you know. In this case, it was useful to remember that \[1 - \cos^2(x) = \sin^2(x)\] Therefore prompting one to multiply the denominator and numerator by \[1+\cos(x)\] in order to get that.

OpenStudy (anonymous):

thank you...:)

OpenStudy (anonymous):

when x = 0 the function becomes 0 / 0 ie indeterminate so it might be a good idea to use l'hopitals rule so differentiate top and bottom

OpenStudy (anonymous):

i haven't studies L Hospital rule yet.. i dont know that..:(

OpenStudy (anonymous):

f'(x) = ( x sec^2 x + tan x) / sin x

OpenStudy (anonymous):

oh ok

OpenStudy (jamesj):

Actually, it's not necessary to apply l'Hopital's rule. you have x(1 + cos x)tan x / sin^2 x. Now as tan x = sin x / cos x, you can cancel one term of sin x in numerator and denominator. Now bring the x into the denominator and you have \[ \frac{(1 + \cos x)\cos x }{(\sin x)/x} \] Now the limit as x --> 0 of sin x / x is 1. Hence you should now now be able to evaluate the limit.

OpenStudy (anonymous):

so answer is 2?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

can we do lim (tan 7x-3x/7x-sin^2 x) x->0 in the same way?

OpenStudy (anonymous):

please tell me what all identities should i use?

OpenStudy (anonymous):

Could you write that using the "Equation" functionality here? It makes it easier to read. Right now it's confusing as there could be parenthesis missing.

OpenStudy (anonymous):

okie..

OpenStudy (anonymous):

\[\left[\begin{matrix}\tan 7x & -3x \\ 7x- & \sin^2x\end{matrix}\right]\]

OpenStudy (anonymous):

Uhm... xD

OpenStudy (anonymous):

Don't think that's what you meant.

OpenStudy (anonymous):

there is a division in the middle.i coudnt get it..

OpenStudy (anonymous):

\[(\tan(7x)-3x)/(7x-\sin^2(x))\] is that it?

OpenStudy (anonymous):